Demand is P=12-(Q/2) and supply is P=Q/4, what is the equilibrium price and quantity?
1
Expert's answer
2016-11-24T12:39:19-0500
To find the equilibrium use the demand and supply curves:
P=12-(Q/2)
P=Q/4
12-(Q/2)= Q/4

Q=16

When Q is equal to 16 units of corn, the price of each unit of corn is $4.